OpenStudy (daenio):

A spring with a force constant of 167 N/m is compressed 1.0 cm. The work done on the spring is ___.

OpenStudy (anonymous):

here's what i think: (do correct me if i'm wrong!) the work done in compressing the spring will be stored as the elastic potential energy of the spring, right? so W = 1/2 kx^2. values of both k and x are given. find W.

OpenStudy (dominusscholae):

Correct! Just make sure you you convert everything to the appropriate units.

OpenStudy (daenio):

I thought it was like this... 1.0 cm / 100 = 0.01m Fs= kx Fs= (167 N/m) (0.01m) Fs= 1.67 N

OpenStudy (vincent-lyon.fr):

What you have calculated above is the force exerted when spring is compressed. It is not the work done to compress it.

OpenStudy (daenio):

Es= 1/2kx2 = ½ 167 (0.01)^2 = 0.00835 J = 8.4 x 10^-3 J

OpenStudy (daenio):

How about that? @Vincent-Lyon.Fr

OpenStudy (anonymous):

@Daenio yup. that's the right answer.

OpenStudy (daenio):

Alright, thanks!